OpenStudy (anonymous):

what is the equation of a line passing through (2,-1) and parallel to the line represented by the equation y=2x+1?

OpenStudy (anonymous):

The required line is parallel to the given line, meaning that their slopes must be equal. slope of the given line = 2 (by comparing y = mx + c) Therefore, slope of the required line must also be = 2 and it must pass through the given point. Therefore, using the slope-point form of the line equation, we can find out the equation. y - y1 = m(x - x1) Therefore y - (-1) = 2 * (x-2) y+1 = 2x - 4 y = 2x - 5 is the required answer

OpenStudy (anonymous):

Got it. Thank You so much!

OpenStudy (anonymous):

What is the equation of a line that is parallel to the line whose equation is y=X+2?